# Self-Improving Robust Preference Optimiziation

### v1

@@ Abstract
Figure 1: SRPO vs. human win rates (WR). We demonstrate robustness by training on TL;DR and evaluating on XSum. Gains on DPO are reported in text captions.
@@ 1 Introduction
(Step 2) Robust Preference Optimization of Generative Model: The next step is to exploit the self-improvement policy learned in the previous step to learn a generative LLM, $\pi$. The key idea here is that the best generative policy can be identified as a policy that generates completions requiring minimal improvement using the optimal self-improvement policy $\pi_{\dagger}$ derived in step $1$. This goal can be achieved by minimizing the objective of step $1$ with respect to the generative policy for in-context completions, $y$. Similar to step $1$, this problem, in its KL-regularized form, can also be solved analytically in terms of the optimal improvement policy $\pi_{\dagger}$ and the optimal generative policy $\pi$. More significantly, we show that the solution for steps $1$ and $2$ can be estimated jointly through a single supervised direct preference optimization scheme using only a dataset of annotated pair-wise completions. Thus, one can solve both for the self-improvement policy $\pi_{\dagger}$ and $\pi$ by minimizing the supervised learning objective of SRPO. Unlike existing RLHF methods, this solution is independent of the behavior policy and is therefore robust to its changes.
As using the self-improvement model in SRPO is a significant departure from the existing paradigm for RLHF, we provide a high-level motivation for it in Sec. 2. We then formalize our objective for SRPO in Sec. 3, allowing for the joint optimization of both $\pi_{\dagger}$ and $\pi$ by optimizing an adversarial min-max objective. In Sec. 4 we present our main algorithmic/mathematical contribution: we prove that the preference probability $p$ can be expressed in terms of the log-likelihoods of the optimal self-improvement policy $\pi_{\dagger}^{*}$ and the log-likelihoods of the optimal robust generative policy $\pi^{*}$. This theoretical finding is the key result for SRPO: solving this system of equations through least-squares regression provides us with the practical supervised SRPO objective that solves for both policy and robust generative policy through a single supervised objective without any need for reward model or online inference. Our key theoretical finding is similar to the main result of DPO (Rafailov et al., 2023) in that both express preference probabilities in terms of the optimal policy. However, DPO result only holds when preference probabilities conform to Bradly-Terry model (Bradley and Terry, 1952), whereas our key result is general as it holds across all preference models. In Sec. 5 we further illustrate our argument on the robustness of SRPO by providing an in-depth analysis of the solution of SRPO and other direct preference optimization methods. We also showcase/analyze the robustness of SRPO on a simple synthetic example. Finally in Sec. 7 we conduct large-scale experiments on training LLMs with SRPO both on in-distribution and OOD summarization tasks and we compare the results with those of standard baselines.
@@ 3 SRPO Objective
with the KL-regularization terms are defined as $D_{\text{KL}}(\pi_{\dagger}||\pi_{\text{ref}}|y_{1},x)=\text{KL}(\pi_{\dagger}(\cdot|y_{1},x)||\pi_{\text{ref}}(\cdot|y_{1},x))$ and $D_{\text{KL}}(\pi||\pi_{\text{ref}}|x)=\text{KL}(\pi(\cdot|x)||\pi_{\text{ref}}(\cdot|x))$.
@@ 6 Related works
Offline preference optimization was introduce as an alternative to more classic RLHF approaches, such as PPO (Schulman et al., 2017; Ouyang et al., 2022) or more generally policy-gradient-based approaches (Roit et al., 2023; Ahmadian et al., 2024). These methods require training a reward model on a preference dataset, usually with a Bradley-Terry model (Bradley and Terry, 1952). The reward model is then used to fine-tune the LLM via online RL, requiring many generations from the model. This reward model shares the common issue of DPO and other direct preference alignment methods, it is dependent on the sampling distribution $\mu$ used for constructing the preference dataset, contrary to SRPO. Moreover, classic RLHF is online, while SRPO is offline and thus more easily scalable.
@@ 7 Experiments
Supervised-fine Tuning. In the SFT stage, we train for 2 epochs, using the AdamW optimizer (Loshchilov and Hutter, 2019), with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, and $0.1$ weight-decay. We use a cosine decay learning rate (Loshchilov and Hutter, 2017) with a peak value of $2\times 10^{-5}$ and $3\%$ of all steps being warm-up steps. We use an effective batch-size of 64.

### v2

# Self-Improving Robust Preference Optimization

@@ 1 Introduction
(Step 2) Robust Preference Optimization of Generative Model: The next step is to exploit the self-improvement policy learned in the previous step to learn a generative LLM, $\pi$.
The key idea here is that the best generative policy can be identified as a policy that generates completions requiring minimal improvement using the optimal self-improvement policy $\pi_{\dagger}$ derived in step $1$. This goal can be achieved by minimizing the objective of step $1$ with respect to the generative policy for in-context completions, $y$. Similar to step $1$, this problem, in its KL-regularized form, can also be solved analytically in terms of the optimal improvement policy $\pi_{\dagger}$ and the optimal generative policy $\pi$. More significantly, we show that the solution for steps $1$ and $2$ can be estimated jointly through a single supervised direct preference optimization scheme using only a dataset of annotated pair-wise completions. Thus, one can solve both for the self-improvement policy $\pi_{\dagger}$ and $\pi$ by minimizing the supervised learning objective of SRPO. Unlike existing RLHF methods, this solution is independent of the behavior policy and is therefore robust to its changes.
As using the self-improvement model in SRPO is a significant departure from the existing paradigm for RLHF, we provide a high-level motivation for it in Sec. 2. We then formalize our objective for SRPO in Sec. 3, allowing for the joint optimization of both $\pi_{\dagger}$ and $\pi$ by optimizing an adversarial min-max objective. In Sec. 4 we present our main algorithmic/mathematical contribution: we prove that the preference probability $p$ can be expressed in terms of the log-likelihoods of the optimal self-improvement policy $\pi_{\dagger}^{*}$ and the log-likelihoods of the optimal robust generative policy $\pi^{*}$. This theoretical finding is the key result for SRPO: solving this system of equations through least-squares regression provides us with the practical supervised SRPO objective that solves for both policy and robust generative policy through a single supervised objective without any need for reward model or online inference. Our key theoretical finding is similar to the main result of DPO (Rafailov et al., 2023) in that both express preference probabilities in terms of the optimal policy. However, DPO result only holds when preference probabilities conform to Bradly-Terry model (Bradley & Terry, 1952), whereas our key result is general as it holds across all preference models. In Sec. 5 we further illustrate our argument on the robustness of SRPO by providing an in-depth analysis of the solution of SRPO and other direct preference optimization methods. We also showcase/analyze the robustness of SRPO on a simple synthetic example. Finally in Sec. 7 we conduct large-scale experiments on training LLMs with SRPO both on in-distribution and OOD summarization tasks and we compare the results with those of standard baselines.
@@ 3 SRPO Objective
with the KL-regularization terms are defined as: $D_{\text{KL}}(\pi_{\dagger}||\pi_{\text{ref}}|y_{1},x)=\text{KL}(\pi_{\dagger}(\cdot|y_{1},x)||\pi_{\text{ref}}(\cdot|y_{1},x))$ and $D_{\text{KL}}(\pi||\pi_{\text{ref}}|x)=\text{KL}(\pi(\cdot|x)||\pi_{\text{ref}}(\cdot|x))$.
@@ 6 Related works
Offline preference optimization was introduce as an alternative to more classic RLHF approaches, such as PPO (Schulman et al., 2017; Ouyang et al., 2022) or more generally policy-gradient-based approaches (Roit et al., 2023; Ahmadian et al., 2024). These methods require training a reward model on a preference dataset, usually with a Bradley-Terry model (Bradley & Terry, 1952). The reward model is then used to fine-tune the LLM via online RL, requiring many generations from the model. This reward model shares the common issue of DPO and other direct preference alignment methods, it is dependent on the sampling distribution $\mu$ used for constructing the preference dataset, contrary to SRPO. Moreover, classic RLHF is online, while SRPO is offline and thus more easily scalable.
@@ 7 Experiments
Supervised-fine Tuning. In the SFT stage, we train for 2 epochs, using the AdamW optimizer (Loshchilov & Hutter, 2019), with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, and $0.1$ weight-decay. We use a cosine decay learning rate (Loshchilov & Hutter, 2017) with a peak value of $2\times 10^{-5}$ and $3\%$ of all steps being warm-up steps. We use an effective batch-size of 64.